# Computational Complexity of Sufficiency in Decision Problems

### v2

@@ 6.2 Computational Rationality of Over-Modeling
Corollary 6.4 explains the observed absence of “config cleanup” tools in software engineering practice. Engineers who include extra parameters are not exhibiting poor discipline—they are adapting to computational constraints. The problem is not lack of tooling effort; it is mathematical intractability.
6.3 Connection to Observed Practice
These theorems provide mathematical grounding for three widespread engineering behaviors:
1. Configuration files grow over time. Removing parameters requires solving coNP-complete problems. Engineers rationally choose linear maintenance cost over exponential minimization cost.
2. Heuristic model selection dominates. ML practitioners use AIC, BIC, cross-validation instead of optimal feature selection because optimal selection is intractable (Theorem 6.3).
3. “Include everything” is a legitimate strategy. When determining relevance/minimal sufficiency is exponential in the worst case (e.g., $O(2^{n})$ by naive subset enumeration, and under ETH a $2^{\Omega(n)}$ worst-case lower bound in the succinct encoding; see Theorem 4.1), including all $n$ parameters costs $O(n)$. For large $n$, this is the rational choice.
These behaviors are not ad hoc workarounds; under the stated computational model they are rational responses to worst-case intractability. The complexity results provide a mathematical lens: over-modeling is not a failure—it is the rational strategy under the model.
7 Implications for Software Architecture
This section states corollaries for software architecture. The complexity results have direct implications for software engineering practice.
7.1 Why Over-Specification Is Rational
Software architects routinely specify more configuration parameters than strictly necessary. Our results show this is computationally rational:
Corollary 7.1 (Rational Over-Specification).
Given a software system with $n$ configuration parameters, checking whether a proposed subset suffices is coNP-complete. Finding the minimum such set is also coNP-complete.
This explains why configuration files grow over time: removing “unnecessary” parameters requires solving a hard problem.
@@ 7.2 Architectural Decision Quotient
The sufficiency framework suggests a measure for architectural decisions:
@@
High ADQ means the configuration subset $I$ leaves many behaviors achievable—it doesn’t constrain the system much. Low ADQ means $I$ strongly constrains behavior.
Proposition 7.3.
Decisions with low ADQ (strongly constraining) require fewer additional decisions to fully specify system behavior.
7.3 Practical Recommendations
Based on our theoretical results:
  1. 1.
    Accept over-modeling: Don’t penalize engineers for including “extra” parameters. The alternative (minimal modeling) is computationally hard.
  2. 2.
    Use bounded scenarios: When the scenario space is small (Proposition 2.10), minimal modeling becomes tractable.
  3. 3.
    Exploit structure: Tree-structured dependencies, bounded alternatives, and separable utilities admit efficient algorithms.
  4. 4.
    Invest in heuristics: For general problems, develop domain-specific heuristics rather than seeking optimal solutions.
@@ 7.4 Hardness Distribution: Right Place vs Wrong Place
A general principle emerges from the complexity results: problem hardness is conserved and is distributed across a system in qualitatively different ways.
@@
Let $P$ be a problem with intrinsic hardness $H(P)$ (measured in computational steps, implementation effort, or error probability). A solution architecture $S$ partitions this hardness into:
@@
Definition 7.5 (Hardness Conservation Principle).
For any problem $P$ with intrinsic hardness $H(P)$, any solution $S$ satisfies:
$H_{\text{central}}(S)+H_{\text{distributed}}(S)\geq H(P).$
This statement is presented as a definitional principle grounded in the prior definition of intrinsic hardness: any correct solution must account for at least $H(P)$ units of work, which the architecture may allocate centrally or distribute across use sites. The substantive, theorem-like consequences (for example, Centralization Dominance) follow from this grounding.
@@
High $\eta$ indicates centralized hardness (paid once); low $\eta$ indicates distributed hardness (paid repeatedly).
@@
For $n>1$ use sites, solutions with higher $H_{\text{central}}$ and lower $H_{\text{distributed}}$ yield:
@@
    Lower total realized hardness: $H_{\text{total}}(S_{1})<H_{\text{total}}(S_{2})$ when $H_{\text{distributed}}(S_{1})<H_{\text{distributed}}(S_{2})$
@@
    Higher leverage: one unit of central effort affects $n$ sites
@@
(1) follows from the total hardness formula. (2) follows from error site counting. (3) follows from the definition of leverage as $L=\Delta\text{Effect}/\Delta\text{Effort}$.  $\blacksquare$
Corollary 7.8 (Right Hardness vs Wrong Hardness).
A solution exhibits hardness in the right place when:
  • Hardness is centralized (high $H_{\text{central}}$, low $H_{\text{distributed}}$)
  • Hardness is paid at design/compile time rather than runtime
  • Hardness is enforced by tooling (type checker, compiler) rather than convention
A solution exhibits hardness in the wrong place when:
  • Hardness is distributed (low $H_{\text{central}}$, high $H_{\text{distributed}}$)
  • Hardness is paid repeatedly at each use site
  • Hardness relies on human discipline rather than mechanical enforcement
Example: Type System Instantiation. Consider a capability $C$ (e.g., provenance tracking) that requires hardness $H(C)$:
@@
For $n$ use sites, manual implementation costs $n\cdot H_{\text{distributed}}$, growing without bound. Native support costs $H_{\text{central}}$ once, amortized across all uses. The “simpler” approach (manual) is only simpler at $n=1$; for $n>H_{\text{central}}/H_{\text{distributed}}$, native support dominates.
Remark 7.9 (Connection to Decision Quotient).
The decision quotient (Section 2) measures which coordinates are decision-relevant. Hardness distribution measures where the cost of handling those coordinates is paid. A high-axis system makes relevance explicit (central hardness); a low-axis system requires users to track relevance themselves (distributed hardness).
Mechanized strengthening (sketch).
We formalized a strengthened reduction in the mechanization: given a Boolean formula $\varphi$ over $n$ variables we construct a decision problem with $n$ coordinates so that if $\varphi$ is not a tautology then every coordinate is decision-relevant. Intuitively, the construction places a copy of the base gadget at each coordinate and makes the global “accept” condition hold only when every coordinate’s local test succeeds; a single falsifying assignment at one coordinate therefore changes the global optimal set, witnessing that coordinate’s relevance. The mechanized statement and proof appear in the development as Reduction_AllCoords.lean (the lemma all_coords_relevant_of_not_tautology).
@@
8 Corollary: Complexity Conservation
A quantitative consequence of the hardness results: when a model handles fewer dimensions than required, the gap must be paid at each use site.
@@ 8 Corollary: Complexity Conservation
Theorem 8.2 (Conservation).
$|\text{Gap}(M,P)|+|R(P)\cap A(M)|=|R(P)|$. The total cannot be reduced—only redistributed between “handled natively” and “handled externally.”
@@
For $n$ decision sites: $\text{TotalExternalWork}=n\times|\text{Gap}(M,P)|$.
@@
Let $H_{\text{central}}$ be the one-time cost of using a complete model. There exists $n^{*}=H_{\text{central}}/|\text{Gap}|$ such that for $n>n^{*}$, the complete model has lower total cost.
Corollary 8.5.
Since identifying $R(P)$ is coNP-complete (Theorem 3.6), minimizing the expressive gap is also intractable.
These results are machine-checked in Lean 4 (HardnessDistribution.lean).

### v3

@@ 7.2 Cost Asymmetry Under ETH
Corollary 7.6 is a formal boundary statement: an always-correct polynomial-time minimizer for arbitrary succinct inputs would collapse P and coNP.
Remark 7.8 (FPT Scope Caveat).
The practical force of worst-case hardness depends on instance structure, especially $k^{*}$. If SUFFICIENCY-CHECK is FPT in parameter $k^{*}$, then small-$k^{*}$ families can remain tractable even under succinct encodings. The strengthened mechanized gadget (all_coords_relevant_of_not_tautology) still proves existence of hard families with $k^{*}=n$; what is typical in deployed systems is an empirical question outside this formal model.
7.3 Regime-Conditional Operational Corollaries
Theorems 7.4 and 7.5 yield the following conditional operational consequences:
1. Conservative retention under unresolved relevance. If irrelevance cannot be certified efficiently under the active regime, the admissible policy is to retain a conservative superset of parameters.[C:T7.4;R:S]
2. Heuristic selection as weakened-guarantee mode. Under [S+ETH], exact global minimization can be exponentially costly in the worst case (Theorem 7.5); methods such as AIC/BIC/cross-validation instantiate the “weaken guarantees” branch of Definition 3.9 [1, 56, 66, 25, 50].[C:T7.5;R:S+ETH]
3. Full-parameter inclusion as an $O(n)$ upper-bound strategy. Under [S+ETH], if exact minimization is unresolved, including all $n$ parameters is an $O(n)$ maintenance upper-bound policy that avoids false irrelevance claims.[C:T7.5;R:S+ETH]
4. Irrationality outside attempted-competence-failure conditions. If the active regime admits exact competence (tractable structural-access conditions or exact relevance identifiability), or if exact competence was never actually attempted, continued over-specification is irrational; hardness no longer justifies it for the stated objective.[C:T6.1;P3.60;R:TA] (L: CC71, IC6, S2P1.)
These corollaries are direct consequences of the hardness/tractability landscape: over-specification is an attempted-competence-failure policy, not a default optimum. Outside that cell, the admissible exits are to shift to tractable regimes from Theorem 6.1 or adopt explicit approximation commitments.[C:T6.1;P3.60;R:RG]
8 Dominance Theorems for Hardness Placement
Regime for this section: the mechanized Boolean-coordinate model [S_bool] plus the architecture cost model defined below. The centralization-vs-distribution framing follows established software-architecture criteria about modular decomposition and long-run maintenance tradeoffs [49, 10, 6, 60]. Although phrased in software terms, the formal objects are substrate-agnostic: the same hardness-placement and amortization equations apply to physical architectures (for example, centralized vs distributed sensing/actuation computation) once mapped into the same coordinate/cost model. [C:P2.34;T2.37,8.16,9.3,9.6;R:AR,LC] (L: CT6,CT4,HD1,HD4,HD3.)
8.1 Over-Specification as Diagnostic Signal
Corollary 8.1 (Persistent Over-Specification).
[C:C8.1;R:H=succinct-hard] In the mechanized Boolean-coordinate model, if a coordinate is relevant and omitted from a candidate set $I$, then $I$ is not sufficient. (L: S2P6)
Proof.
This is the contrapositive of S2P3.  $\blacksquare$
Corollary 8.2 (Exact Relevance Identifiability Criterion).
[C:C8.2;R:H=succinct-hard] In the mechanized Boolean-coordinate model, for any candidate set $I$:
$I\text{ is exactly relevance-identifying}\iff\bigl(I\text{ is sufficient and }I\subseteq R_{\mathrm{rel}}\bigr),$
where $R_{\mathrm{rel}}$ is the full relevant-coordinate set. (L: S2P1)
Proof.
This is exactly S2P1, with $R_{\mathrm{rel}}=\texttt{relevantFinset}$.  $\blacksquare$
Remark 8.3 (Over-Specification Is Regime-Conditional, Not Default).
In this paper’s formal typing, persistent over-specification is admissible only under attempted competence failure: after an attempted exact procedure, exact relevance competence remains unavailable in the active regime, so integrity retains uncertified coordinates (Section 3; Section 7).[C:P3.60;R:AR] Once exact competence is available in the active regime (Corollaries 8.78.9 together with Corollary 8.2), certifiably irrelevant coordinates are removable; persistent over-specification is irrational for the same objective (verified total-cost minimization).[C:T6.1;P3.60;R:TA] (L: IC18, IC25, IC7, L15, L16, L17, S2P1.)
@@ 8.2 Architectural Decision Quotient
Proposition 8.5 (ADQ Ordering).
For coordinate sets $I,J$ in the same system, if $\mathrm{ADQ}(I)<\mathrm{ADQ}(J)$, then fixing $I$ leaves a strictly smaller achievable-behavior set than fixing $J$. (L: CC2)
Proof.
The denominator $|B|$ is shared. Thus $\mathrm{ADQ}(I)<\mathrm{ADQ}(J)$ is equivalent to a strict inequality between the corresponding achievable-behavior set cardinalities.  $\blacksquare$
8.3 Corollaries for Practice
Corollary 8.6 (Cardinality Criterion for Exact Minimization).
[C:C8.6;R:H=succinct-hard] In the mechanized Boolean-coordinate model, existence of a sufficient set of size at most $k$ is equivalent to the relevance set having cardinality at most $k$. (L: S2P2)
Proof.
By S2P2, sufficiency of size $\leq k$ is equivalent to a relevance-cardinality bound $\leq k$ in the Boolean-coordinate model.  $\blacksquare$
Corollary 8.7 (Bounded-Regime Tractability).
[C:C8.7;R:H=tractable-structured] When the bounded-action or explicit-state conditions of Theorem 6.1 hold, minimal modeling can be solved in polynomial time in the stated input size. (L: L15)
Proof.
This is the bounded-regime branch of Theorem 6.1, mechanized as L3.  $\blacksquare$
Corollary 8.8 (Separable-Utility Tractability).
[C:C8.8;R:H=tractable-structured] When utility is separable with explicit factors, sufficiency checking is polynomial in the explicit-state regime. (L: L16)
Proof.
This is the separable-utility branch of Theorem 6.1, mechanized as L4.  $\blacksquare$
Corollary 8.9 (Tree-Structured Tractability).
[C:C8.9;R:H=tractable-structured] When utility factors form a tree structure with explicit local factors, sufficiency checking is polynomial in the explicit-state regime. (L: L17)
Proof.
This is the tree-factor branch of Theorem 6.1, mechanized as L5.  $\blacksquare$
Corollary 8.10 (Mechanized Hard Family).
There is a machine-checked family of reduction instances where, for non-tautological source formulas, every coordinate is relevant ($k^{*}=n$), exhibiting worst-case boundary complexity. (L: L1)
Proof.
The strengthened reduction proves that non-tautological source formulas induce instances where every coordinate is relevant; this is mechanized as L1.  $\blacksquare$
@@ 8.4 Hardness Distribution: Right Place vs Wrong Place
Let $P$ be a problem family under the succinct encoding of Section 2.4. In this section, baseline hardness $H(P;n)$ denotes worst-case computational step complexity on instances with $n$ coordinates (equivalently, as a function of succinct input length $L$) in the fixed encoding regime. A solution architecture $S$ partitions this baseline hardness into:
@@
Proposition 8.12 (Baseline Lower-Bound Principle).
[C:P8.12;R:H=cost-growth] For any problem family $P$ measured by $H(P;n)$ above, any solution architecture $S$ and any number of use sites $n\geq 1$, if $H_{\text{total}}(S)$ is measured in the same worst-case step units over the same input family, then:
$H_{\text{total}}(S)=H_{\text{central}}(S)+n\cdot H_{\text{distributed}}(S)\geq H(P;n).$
For SUFFICIENCY-CHECK, Theorem 5.1 provides the baseline on the hard succinct family: $H(\text{SUFFICIENCY-CHECK};n)=2^{\Omega(n)}$ under ETH. (L: HD20)
Proof.
By definition, $H(P;n)$ is a worst-case lower bound for correct solutions in this encoding regime and cost metric. Any such solution architecture decomposes total realized work as $H_{\text{central}}+n\cdot H_{\text{distributed}}$, so that total cannot fall below the baseline.  $\blacksquare$
@@
(L: HD10)
Proposition 8.14 (Efficiency Equivalence).
[C:P8.14;R:H=cost-growth] For fixed $n$ and positive total hardness, larger $\eta(S,n)$ is equivalent to a larger central share of realized hardness. (L: HD10)
Proof.
From Definition 8.13, $\eta(S,n)$ is exactly the fraction of total realized hardness paid centrally.  $\blacksquare$
Definition 8.15 (Right vs Wrong Hardness Placement).
For a solution architecture $S$ in this linear model:
$\text{right hardness}\iff H_{\mathrm{distributed}}(S)=0,\qquad\text{wrong hardness}\iff H_{\mathrm{distributed}}(S)>0.$
(L: HD11, HD12)
@@
Let $S_{\mathrm{right}},S_{\mathrm{wrong}}$ be architectures over the same problem family with
$H_{\mathrm{distributed}}(S_{\mathrm{right}})=0,\quad H_{\mathrm{central}}(S_{\mathrm{right}})>0,\quad H_{\mathrm{distributed}}(S_{\mathrm{wrong}})>0,$
and let $n>\max\!\bigl(1,H_{\mathrm{central}}(S_{\mathrm{right}})\bigr)$. Then:
@@
    Lower total realized hardness:
    $H_{\text{total}}(S_{\mathrm{right}})<H_{\text{total}}(S_{\mathrm{wrong}})$
@@
    Quantified leverage: moving one unit of work from distributed to central saves exactly n1n-1 units of total realized hardness
(L: HD1, HD2)
@@
(1) and (2) are exactly the bundled theorem HD1. (3) is exactly HD2.  $\blacksquare$
Corollary 8.17 (Right-Place vs Wrong-Place Hardness).
[C:C8.17;R:H=cost-growth] In the linear model, a right-hardness architecture strictly dominates a wrong-hardness architecture once use-site count exceeds central one-time hardness. Formally, for architectures $S_{\mathrm{right}},S_{\mathrm{wrong}}$ over the same problem family, if $S_{\mathrm{right}}$ has right hardness, $S_{\mathrm{wrong}}$ has wrong hardness, and $n>H_{\mathrm{central}}(S_{\mathrm{right}})$, then
$H_{\mathrm{central}}(S_{\mathrm{right}})+n\,H_{\mathrm{distributed}}(S_{\mathrm{right}})<H_{\mathrm{central}}(S_{\mathrm{wrong}})+n\,H_{\mathrm{distributed}}(S_{\mathrm{wrong}}).$
(L: HD16)
Proof.
This is the mechanized theorem HD10.  $\blacksquare$
Proposition 8.18 (Dominance Modes).
[C:P8.18;R:H=cost-growth] This section uses two linear-model dominance modes plus generalized nonlinear dominance and boundary modes:
  1. 1.
    Strict threshold dominance: Corollary 8.17 gives strict inequality once $n>H_{\mathrm{central}}(S_{\mathrm{right}})$.
  2. 2.
    Global weak dominance: under the decomposition identity used in HD3, centralized hardness placement is never worse for all $n\geq 1$.
  3. 3.
    Generalized nonlinear dominance: under bounded-vs-growing site-cost assumptions (Theorem 8.22), right placement strictly dominates beyond a finite threshold without assuming linear per-site cost.
  4. 4.
    Generalized boundary mode: without those growth-separation assumptions, strict dominance is not guaranteed (Proposition 8.24).
Proof.
Part (1) is Corollary 8.17. Part (2) is exactly HD3. Part (3) is Theorem 8.22. Part (4) is Proposition 8.24.  $\blacksquare$
Illustrative Instantiation (Type Systems). Consider a capability $C$ (e.g., provenance tracking) with one-time central cost $H_{\text{central}}$ and per-site manual cost $H_{\text{distributed}}$:
@@
The table is schematic; the formal statement is Corollary 8.19. The type-theoretic intuition behind this instantiation is consistent with classic accounts of abstraction and static interface guarantees [12, 53, 51, 36, 62, 22].
Corollary 8.19 (Type-System Threshold).
[C:C8.19;R:H=cost-growth] For the formal native-vs-manual architecture instance, native support has lower total realized cost for all
$n>H_{\mathrm{baseline}}(P),$
where $H_{\mathrm{baseline}}(P)$ corresponds to the Lean identifier intrinsicDOF(P) in module HardnessDistribution. (L: HD14)
Proof.
Immediate from HD9.  $\blacksquare$
8.5 Extension: Non-Additive Site-Cost Models
Definition 8.20 (Generalized Site Accumulation).
Let $C_{S}:\mathbb{N}\to\mathbb{N}$ be a per-site accumulation function for architecture $S$. Define generalized total realized hardness by
$H_{\text{total}}^{\mathrm{gen}}(S,n)=H_{\text{central}}(S)+C_{S}(n).$
Definition 8.21 (Eventual Saturation).
A cost function $f:\mathbb{N}\to\mathbb{N}$ is eventually saturating if there exists $N$ such that for all $n\geq N$, $f(n)=f(N)$.
Theorem 8.22 (Generalized Dominance by Growth Separation).
[C:T8.22;R:H=cost-growth] Let
$H_{\text{total}}^{\mathrm{gen}}(S,n)=H_{\text{central}}(S)+C_{S}(n).$
For two architectures $S_{\mathrm{right}},S_{\mathrm{wrong}}$, suppose there exists $B\in\mathbb{N}$ such that:
  1. 1.
    $C_{S_{\mathrm{right}}}(m)\leq B$ for all $m$ (bounded right-side per-site accumulation),
  2. 2.
    $m\leq C_{S_{\mathrm{wrong}}}(m)$ for all $m$ (identity-lower-bounded wrong-side growth).
Then for every
$n>H_{\text{central}}(S_{\mathrm{right}})+B,$
one has
$H_{\text{total}}^{\mathrm{gen}}(S_{\mathrm{right}},n)<H_{\text{total}}^{\mathrm{gen}}(S_{\mathrm{wrong}},n).$
(L: HD8)
Proof.
This is exactly the mechanized theorem HD6.  $\blacksquare$
Corollary 8.23 (Eventual Generalized Dominance).
[C:C8.23;R:H=cost-growth] If condition (1) above holds and there exists $N$ such that condition (2) holds for all $m\geq N$, then there exists $N_{0}$ such that for all $n\geq N_{0}$,
$H_{\text{total}}^{\mathrm{gen}}(S_{\mathrm{right}},n)<H_{\text{total}}^{\mathrm{gen}}(S_{\mathrm{wrong}},n).$
(L: HD9)
Proof.
Immediate from HD7.  $\blacksquare$
Proposition 8.24 (Generalized Assumption Boundary).
[C:P8.24;R:H=cost-growth] In the generalized model, strict right-vs-wrong dominance is not unconditional. There are explicit counterexamples:
  1. 1.
    If wrong-side growth lower bounds are dropped, right-side strict dominance can fail for all $n$.
  2. 2.
    If right-side boundedness is dropped, strict dominance can fail for all $n$ even when wrong-side growth is linear.
(L: HD7, HD6)
Proof.
This is exactly the pair of mechanized boundary theorems listed above.  $\blacksquare$
Theorem 8.25 (Linear Model: Saturation iff Zero Distributed Hardness).
[C:T8.25;R:H=cost-growth] In the linear model of this section,
$H_{\text{total}}(S,n)=H_{\text{central}}(S)+n\cdot H_{\text{distributed}}(S),$
the function $n\mapsto H_{\text{total}}(S,n)$ is eventually saturating if and only if $H_{\text{distributed}}(S)=0$. (L: HD19)
Proof.
This is exactly the mechanized equivalence theorem above.  $\blacksquare$
Theorem 8.26 (Generalized Model: Saturation is Possible).
[C:T8.26;R:H=cost-growth] There exists a generalized site-cost model with eventual saturation. In particular, for
$C_{K}(n)=\begin{cases}n,&n\leq K\\ K,&n>K,\end{cases}$
both $C_{K}$ and $n\mapsto H_{\text{central}}+C_{K}(n)$ are eventually saturating. (L: HD17, HD5)
Proof.
This is the explicit construction mechanized in Lean.  $\blacksquare$
Corollary 8.27 (Positive Linear Slope Cannot Represent Saturation).
[C:C8.27;R:H=cost-growth] No positive-slope linear per-site model can represent the saturating family above for all $n$. (L: HD15)
Proof.
This follows from the mechanized theorem that any linear representation of the saturating family must have zero slope.  $\blacksquare$
Mechanized strengthening reference.
The strengthened all-coordinates-relevant reduction is presented in Section 4 (“Mechanized strengthening”) and formalized in Reduction_AllCoords.lean via all_coords_relevant_of_not_tautology.
@@ 8.5 Extension: Non-Additive Site-Cost Models
9 Conservation Law for Decision-Relevant Coordinates
The load-bearing fact in this section is not the set identity itself; it is the difficulty of shrinking the required set $R(P)$ in the first place. By Theorem 4.6 (and Theorem 4.7 for minimization), exact relevance identification is intractable in the worst case under succinct encoding. The identities below therefore quantify how unresolved relevance is redistributed between central and per-site work.[C:T4.6,4.7;R:S] For encoded physical systems, the same redistribution laws apply under declared physical-to-core transport: unresolved coordinates remain per-site external work, while resolved coordinates are paid once in centralized structure.[C:P2.34;T9.3,9.5,9.6;R:AR,LC] (L: CT6,HD4,HD21,HD3.)
@@ 9 Conservation Law for Decision-Relevant Coordinates
Definition 9.2 (Simplicity Tax).
The simplicity tax is the size of the expressive gap:
$\text{SimplicityTax}(M,P):=|\text{Gap}(M,P)|.$
Theorem 9.3 (Redistribution Identity).
$|\text{Gap}(M,P)|+|R(P)\cap A(M)|=|R(P)|$. The total cannot be reduced, only redistributed between “handled natively” and “handled externally.” (L: HD4)
Proof.
In the finite-coordinate model this is the exact set-cardinality identity
$|R\setminus A|+|R\cap A|=|R|,$
formalized as HD5.  $\blacksquare$
Remark 9.4 (Why this is nontrivial in context).
The algebraic identity in Theorem 9.3 is elementary. Its force comes from upstream hardness: reducing $|R(P)|$ by exact relevance minimization is worst-case intractable under the succinct encoding, so redistribution is often the only tractable lever available.
@@
For $n$ decision sites:
$\text{TotalExternalWork}=n\times\text{SimplicityTax}(M,P).$
(L: HD21)
Proof.
This is by definition of per-site externalization and is mechanized as HD12.  $\blacksquare$
@@
Let $H_{\text{central}}$ be the one-time cost of using a complete model. There exists
$n^{*}=\left\lfloor\frac{H_{\text{central}}}{\text{SimplicityTax}(M,P)}\right\rfloor$
such that for $n>n^{*}$, the complete model has lower total cost. (L: HD3)
Proof.
For positive per-site tax, the threshold inequality
$n>\left\lfloor\frac{H_{\text{central}}}{\text{SimplicityTax}}\right\rfloor\implies H_{\text{central}}<n\cdot\text{SimplicityTax}$
is mechanized as HD4.  $\blacksquare$
Corollary 9.7 (Gap Externalization).
[C:C9.7;R:H=cost-growth] If $\text{Gap}(M,P)\neq\emptyset$, then external handling cost scales linearly with the number of decision sites. (L: HD21, HD18)
Proof.
The exact linear form is HD12. When the gap is nonempty (positive tax), monotone growth with $n$ is HD11.  $\blacksquare$
Corollary 9.8 (Exact Minimization Criterion).
[C:C9.8;R:H=succinct-hard] For mechanized Boolean-coordinate instances, “there exists a sufficient set of size at most kk” is equivalent to “the relevant-coordinate set has cardinality at most kk.” (L: S2P2)
Proof.
This is S2P2.  $\blacksquare$
Appendix A provides theorem statements and module paths for the corresponding Lean formalization.